OpenStudy (anonymous):

Explain in complete sentences and demonstrate how to add:

OpenStudy (anonymous):

\[\frac{ x-5 }{ 2x } \frac{ x+1 }{ 2x }\]

OpenStudy (anonymous):

the two fractions are over each other @SolomonZelman

OpenStudy (solomonzelman):

when you add fractions with the same denominator, only the numerator changes while the denominator remains the same. Add like terms in the numerator and you will get two times a number minus four, which is going to be over two times the number. IDK, ain't good at explaining math using sentences.

OpenStudy (anonymous):

could you demonstrate?

OpenStudy (solomonzelman):

\[\frac{x-5}{2x}+\frac{x+1}{2x}~~~~~->\]\[\frac{(x-5)+(x+1)}{2x}\]\[\frac{2x+4}{2x}\]

OpenStudy (anonymous):

ohhhh i get it so the denominator doesn't change.

OpenStudy (solomonzelman):

then reduce it,\[\frac{2x+4}{2x}~~~->~~~\frac{2(x+2)}{2x}~~~_>~~~\frac{x+2}{x}\] And yes the denominator stays the same.

OpenStudy (anonymous):

ok i see hold on ill brb

OpenStudy (solomonzelman):

K

OpenStudy (anonymous):

k back

OpenStudy (solomonzelman):

I"m here for a couple minutes or so....

OpenStudy (anonymous):

ok so thats how you do it

OpenStudy (anonymous):

Perform the subtraction: [(x-5)/2x] - [x/(x+3)] Discuss and identify any possible restrictions that exist in the resulting rational expression.

OpenStudy (anonymous):

\[\frac{ x-5 }{ 2x }\]

OpenStudy (anonymous):

minus

OpenStudy (anonymous):

\[\frac{ x }{ x+3 }\]

OpenStudy (solomonzelman):

demonstrate or sentences?

OpenStudy (anonymous):

both, like the first question

OpenStudy (solomonzelman):

I am not good at wrting sentences, but I can show you the math....

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

\[\frac{x-5}{2x}~-~\frac{x}{x+3}\]\[\frac{x \color{red} {\times (x+3)}}{2x\color{red} {\times (x+3)}}~-~\frac{x \color{red} {\times 2x}}{(x+3) \color{red} {\times 2x} }\]

OpenStudy (solomonzelman):

\[\frac{(x^2+3x)}{(2x^2+3x)}~-~\frac{(2x^2)}{(2x^2+3x)}\]

OpenStudy (solomonzelman):

\[\frac{(x^2+3x)-(2x^2)}{(2x^2+3x)}\]

OpenStudy (solomonzelman):

\[\frac{(-x^2+3x)}{(2x^2+3x)}\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

so you cross multiply?

OpenStudy (solomonzelman):

no, I find the common denominator, doing it just like with any normal fraction, except that I am finding the common denominator between binomials and variables.

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

any possible restrictions that are in the resulting expression?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

the denominator cannot equal zero, so \[\color{blue} {2x^2+3x≠0}\]\[\color{blue} {2x^2+3x\color{red} {+0}≠0}\]plug it inside the quadratic formula. \[\color{blue} {\frac{-3±\sqrt{3^2+(4 \times 2 \times 0)}}{2 \times 2}}\]\[\color{blue} {\frac{-3±\sqrt{9}}{4}}\]\[\color{blue} {\frac{-3±3}{4}}\]\[\color{blue} {x≠ 0~~~~~~~~~~or~~~~~~~~x≠ -1\frac{1}{2}}\]

OpenStudy (anonymous):

thanks.